Opened 10 years ago

Closed 2 years ago

#272 closed flight processing (fixed)

EU09/06, flight day 224a/2009, Vatna_V1

Reported by: mark1 Owned by: asm
Priority: alpha 4 medium Milestone: 2009 Data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/other/EU09_06-2009_224a_Vatna_V1

Data arrived from ARSF via SATA disk LETTER L on 7th October 2009.

Scientific objective: Glacier response to a changing climate in southern Iceland

Priority: alpha-4-medium

PI: David Graham

Sensors:

  • Eagle (delivered)
  • Hawk (delivered)
  • Leica LIDAR (delivered)

Change History (14)

comment:1 Changed 10 years ago by mark1

Awaiting logsheet.

comment:2 Changed 9 years ago by chrfi

applanix nav processing completed.

note: supplied base station too far from flight, used basestation downloaded through applanix

note: unable to use data from http://hraun.vedur.is/ja/gps/outdata/glaciermapping/

comment:3 Changed 9 years ago by chrfi

  • Owner set to chrfi
  • Status changed from new to assigned

comment:4 Changed 9 years ago by chrfi

processing hyperspectral

comment:5 Changed 9 years ago by chrfi

added correct prefix to eagle and hawk filenames in arsf-data and workspace

comment:6 Changed 9 years ago by chrfi

sct offsets

line number eagle hawk
00 1.20 0.06
01 -0.07
02 0.03

comment:7 Changed 9 years ago by mark1

Delivery directory for LIDAR created by chrfi at:
~airborne/workspace/EU09_06-2009_224a_Vatna_V1/delivery/20100518b/EU09_06

Checked and now looks OK after corrections. Ready to deliver.

comment:8 Changed 9 years ago by knpa

Delivery checked hyperspectral at:

.../EU09_06-2009_224a_Vatna_V1/delivery/20100518a/

Made some minor edits to ReadMe.
All else looks good apart from the following:

  1. hawk and eagle filenames start at 00, they should start at 01 (remember to update readme too: contents, azgcorr/azexhdf commands)
  2. should be dem readme in dem folder
  3. Readme: no bin dir listed in contents, also dem listing out of alphabetical order

comment:9 Changed 9 years ago by chrfi

waiting for space on bowhead before rsync

comment:10 Changed 8 years ago by knpa

  • Description modified (diff)
  • Owner changed from chrfi to knpa

comment:11 Changed 2 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:12 Changed 2 years ago by asm

Archiving

Archived and it is available from CEDA. Processed data at http://browse.ceda.ac.uk/browse/neodc/arsf/2009/EU09_06/EU09_06-2009_224a_Vatna_V1

Plymouth 12th May 2017

comment:13 Changed 2 years ago by asm

  • Owner changed from knpa to asm
  • Status changed from assigned to new

comment:14 Changed 2 years ago by asm

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.